### v1

# Universal logical operations with a dynamical qubit in Floquet code

Abstract
Quantum error correction (QEC) protects quantum systems against inevitable noises and control inaccuracies, providing a pathway towards fault-tolerant (FT) quantum computation. However, the significant overhead of physical qubits required to encode a single logical qubit poses a major challenge for scalability and practical implementation. Floquet QEC codes, a recent innovation, mitigate this challenge by utilizing time-periodic measurements to introduce additional dynamical logical qubits, thereby enhancing resource efficiency in QEC. Here, we experimentally implement the Floquet-Bacon-Shor code on a superconducting quantum processor. We encode a dynamical logical qubit within a $3\times 3$ lattice of data qubits, alongside a conventional static logical qubit. We demonstrate FT encoding and measurement of the two-qubit logical states and stabilize the states using repeated error detection cycles. Additionally, we showcase universal single-qubit logical gates on the dynamical qubit. By implementing a logical CNOT gate, we entangle the dynamical and static logical qubits, generating an error-detected logical Bell state with a fidelity of 75.9%. Our results highlight the potential of Floquet codes for scalable and resource-efficient FT quantum computation.
I Introduction
Figure 1: Implementation of the FBS code. a, Schematic of the quantum device for the FBS code, consisting of 9 data qubits (white circles labeled as $D_{i}$, where i = 1 to 9) and 12 ancilla qubits (orange circles for $Z$ ancillas and blue circles for $X$ ancillas). The four central qubits are alternately used as $X$ and $Z$ ancillas in different stabilizer rounds. Extra ancilla qubits (grey circles) are used for circuit compression and the construction of logical CNOT gate. Each pair of qubits is connected by a tunable coupler (H shape). b, Illustration of stabilizer measurements for the BS code (left) and the FBS code (right). The FBS code features period-4 stabilizer measurements with evolving logical operators $X_{\rm d}$ (purple lines) and $Z_{\rm d}$ (green lines) for the dynamical qubit in different rounds. In contrast, the logical operators $X_{\rm s}$ and $Z_{\rm s}$ are fixed and identical for both codes. Weight-2 Pauli measurements are used to check the parity of the corresponding data qubits in the $X$ (blue) and $Z$ (orange) bases. c, Cumulative distributions of errors for simultaneous single-qubit gates, CZ gates, ancilla qubit readout, and data qubit idle (with dynamical decoupling) during ancilla readout are depicted. The median error rates are 0.07%, 0.97%, 1.58%, and 1.43%, respectively, as indicated by the vertical lines.
Quantum computers hold potential for solving computational problems that are provably intractable for classical systems, with transformative applications in quantum simulation [1], cryptography [2], and machine learning [3]. However, their practical realization is hindered by a fundamental challenge: quantum information is inherently fragile, susceptible to inevitable environmental noise and control inaccuracies. To address this challenge, quantum error correction (QEC) has emerged as a critical framework [4, 5, 6]. By redundantly encoding logical qubits into larger systems with many physical qubits, QEC detects and corrects errors through repeated syndrome measurements. Combined with fault-tolerant (FT) logical operations, QEC not only safeguards quantum information, but also lays the foundation for scalable and reliable quantum computation.
A variety of QEC codes have been proposed, including surface code [7], color code [8], Bacon-Shor (BS) code [9], and quantum LDPC codes [10, 11]. Significant experimental progresses have been achieved in recent years across various platforms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Notably, exponential error suppression has been demonstrated by scaling up a surface-code logical qubit using superconducting circuits [22]. While high-performance logical qubits can be realized by increasing the size of quantum processors, the code rate–the ratio of logical qubits to physical qubits–remains a crucial metric [11]. Low encoding efficiency can make FT quantum computing resource-intensive, especially for solid-state quantum devices limited to nearest-neighbor coupling.
Recently proposed time-dynamical QEC, including Floquet codes [23, 24] and generalized time-dynamical code implementations [25, 26, 27, 28, 29], opens up new opportunities for fault-tolerant quantum computation. A recent work [30] has shown the flexibility in surface code realization using time-dynamical implementations. Unlike conventional stabilizer or subsystem codes that encode logical information in a fixed subspace, the Floquet codes allow logical information to evolve periodically over time. Additional dynamical logical qubits are encoded through an appropriate measurement schedule, thereby improving the code rate. During the periodic schedule of Pauli measurements, logical subspaces and operators adapt to stabilizer measurements while preserving the encoded information. Despite extensive theoretical study since its inception [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], the Floquet code has not yet been experimentally demonstrated.
In this work, we implement the Floquet-Bacon-Shor (FBS) code [42] on a superconducting quantum processor, encoding one conventional static logical qubit and one dynamical logical qubit simultaneously within a $3\times 3$ array of data qubits. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using multi-round stabilizer measurements. We showcase FT Pauli gates and non-fault-tolerant (nFT) arbitrary rotation gates on the dynamical qubit. Furthermore, we achieve a logical Bell state fidelity of $75.9\%$ by entangling the dynamical and static logical qubits via a logical CNOT gate, with the gate fidelity assessed at $84.1\%$ through logical quantum process tomography (LQPT). With these demonstrated universal two-qubit logical operations, our results highlight the potential of Floquet codes for scalable and resource-efficient QEC.
II Results
Figure 2: Encoding, stabilization, and measurement of two-qubit logical states in the FBS code. a, Simplified circuit for encoding, stabilizing and measuring two-qubit logical states. All physical qubits are initialized to the ground state before the circuit. Period-4 stabilizer measurements are alternately performed following the encoding circuit. b, Encoding circuits for logical Pauli states $|+,0\rangle$ and $|1,0\rangle$. BS stabilizer measurements are included in the latter case, with the dynamical qubit state post-selected to be $|0\rangle$. c, Example logical measurement circuits after stabilizer $S_{Z}^{\rm D}$: FT $ZZ$ measurement (left) and nFT $XZ$ measurement (right). Logical operators are indicated within the dashed-line boxes. Measurement outcomes of some data qubits are used to calculate the values of the $X$ (blue) and $Z$ (orange) stabilizers, enabling error detection. d, LQST fidelities of all 36 two-qubit Pauli states. The 16 states $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}^{\otimes 2}$ are fault-tolerantly prepared and exhibit higher fidelity, as shown to the left of the dashed line. The blue portions indicate the fidelities without error detection, while the green portions highlight the fidelity improvement with error detection in the logical measurement. e, Experimentally measured (filled symbols) and simulated (open symbols) expectation values of $X_{\rm s}Z_{\rm d}$ operator versus stabilizer round, for the encoded states $|-,1\rangle$ (top) and $|-,0\rangle$ (bottom). The solid and dashed lines represent exponential fits. Each stabilizer round lasts 920 ns, comprising 720 ns for readout pulses and resonator photon depletion, and 200 ns for single-qubit and two-qubit gates. The cyan and green colors correspond to data without post-processing and data with measured error syndromes post-selected out, respectively. Due to the exponential decrease in the retained data during error detection, the analysis is limited to 16 rounds, with a total of 100,000 experimental shots. The insets show the error detection probabilities for the weight-6 stabilizers over 24 measurement rounds, with the final round calculated from the logical measurements.
The FBS code we demonstrate is a generalization of the $[[9,1,3]]$ BS code [9], a subsystem code that encodes one logical qubit on a $3\times 3$ square lattice, as illustrated in Fig. 1a. Each vertex represents a data qubit, and each edge corresponds to a weight-2 Pauli measurement. The static logical qubit is encoded in a fixed subsystem, with its logical Pauli operators given by
$X_{\rm s}=X_{4}X_{5}X_{6},\quad Z_{\rm s}=Z_{2}Z_{5}Z_{8},\quad Y_{\rm s}=iX_{\rm s}Z_{\rm s}.$ (1)
The BS code can be stabilized in a single measurement round using only weight-2 Pauli measurements, whose ordering is inconsequential [9], as shown in Fig. 1b. The four weight-6 stabilizers, $S_{X}^{\text{A}}$, $S_{Z}^{\text{B}}$, $S_{X}^{\text{C}}$, and $S_{Z}^{\text{D}}$, are derived by multiplying the outcomes of the corresponding Pauli measurements, enabling the detection and subsequent correction of any single-qubit error. Since these Pauli measurements do not always commute, the Hilbert space of the 9 physical qubits consists of 4 degrees of freedom fixed by the stabilizers, 1 degree for logical information, and 4 additional degrees corresponding to the gauge space, which does not encode logical information. By imposing a period-4 ordering on the weight-2 Pauli measurements, the gauge space dynamically encodes an additional logical qubit, thereby forming the FBS code. In each round, $4$ weight-$2$ Pauli measurements are performed, whose outcomes determine one weight-$6$ stabilizer and also influence the evaluation of dynamical operators (see Methods). The logical Pauli operators $X_{\text{d}}$ and $Z_{\text{d}}$ of the dynamically encoded qubit evolve periodically during the stabilizer measurements, as indicated by the purple and green lines in Fig. 1b. Both the static and dynamical logical qubits are stabilized over the 4 measurement rounds. Although the $3\times 3$ system size is insufficient for error correction of the dynamical qubit, the stabilizers extracted from the measurements are sufficient to detect any single-qubit error. Within the framework of error detection, an operation is considered FT if any single-qubit error produces a non-trivial syndrome and is thereby detectable [47].
We implement the FBS code on a superconducting quantum processor comprising a $6\times 11$ array of Transmon qubits [48], each coupled to its four nearest neighbors through tunable couplers [49]. The experimental setup and device parameters are detailed in Supplementary Information Section II. The static and dynamical logical qubits are encoded on a 21-qubit subset of the processor, consisting of $N_{d}=9$ data qubits and $N_{a}=12$ ancilla qubits. The data qubits, labeled as $D_{i}$ ($i=1$ to 9), form a $3\times 3$ array and are interlaced with $X$ and $Z$ ancilla qubits for stabilizer measurements. Additional ancilla qubits are incorporated to optimize the QEC circuits. High-fidelity quantum gates and quantum non-demolition (QND) readout [50] are implemented on the 21-qubit subset, with the cumulative distributions of errors illustrated in Fig. 1c. The readout fidelities of the ancilla qubits are benchmarked using repeated measurements [51] (see Supplementary Information Section III). Dynamical decoupling (DD) techniques [52] are employed to protect the data qubits from low-frequency noise and cross-measurement-induced dephasing [53, 54, 55]. The idle errors (with DD) on the data qubits during ancilla qubit readout are quantified using interleaved randomized benchmarking [56].
[…]
Logical measurements are performed by measuring the data qubits in their appropriate bases, as illustrated by the example circuits in Fig. 2c. The measurement circuits for the dynamical qubit change according to the logical operator definitions in different stabilizer rounds. The outcome of a logical operator is determined by calculating the total parity of the relevant data qubits (see Methods). Data qubits that do not define the logical operators are also measured to enable error detection. The $Z_{\rm s}Z_{\rm d}$ measurement is FT because the measured outcomes allow for the calculation of stabilizers $S_{Z}^{\rm D}$ and $S_{Z}^{\rm B}$, enabling the detection of any single-qubit error. In contrast, the measurement of $X_{\rm s}Z_{\rm d}$ is nFT; however, the value of stabilizer $S_{X}^{\rm C}$ can be calculated from the measurement outcomes to detect phase-flip errors and measurement errors on qubits $D_{1}$ through $D_{6}$. We present measured fidelities of the 36 encoded states in Fig. 2d. The state fidelities are determined using logical quantum state tomography (LQST), where 9 logical bases $\{Z,X,Y\}^{\otimes 2}$ are measured (see Methods for details on LQST and Supplementary Information Section VI for full measurement circuits). Error detection during the logical measurements improves the state fidelities, as indicated by the green portions in the plot.
Figure 3: Logical gates on the dynamical qubit. a, FT Pauli gates on the dynamical qubit. Transversal Pauli gates are applied to the data qubits and inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as shown at the top. The final state is characterized using LQST. The state infidelities with error detection after applying the Pauli gates on different encoded states are shown below the circuit. For each dynamical qubit state, the infidelity is averaged over 4 static qubit states {$|0\rangle$, $|1\rangle$, |+ket|$\rangle$| + ⟩, |ket|$\rangle$| - ⟩}. Error bars correspond to 95% confidence intervals. b, Logical rotation gates on the dynamical qubit around the $Z_{\rm d}$-axis (left) and $X_{\rm d}$-axis (right). The measurement expectation values of the operators $X_{\rm d}$ (cyan), $Y_{\rm d}$ (green) and $Z_{\rm d}$ (orange) are shown as functions of rotation angles $\varphi$ and $\theta$. The gate circuits (shown at the top) for rotations around the $Z_{\rm d}$-axis ($X_{\rm d}$-axis) are inserted between stabilizers $S_{Z}^{\rm B}$ ($S_{X}^{\rm C}$) and $S_{X}^{\rm C}$ ($S_{Z}^{\rm D}$). The filled (open) symbols represent data with (without) error detection in the four stabilizer rounds and the final logical measurement. The dashed curves correspond to trigonometric fits.
We perform period-4 stabilizer measurements to preserve the encoded two-qubit logical states. Specifically, the four stabilizers $S_{X}^{\rm A}$, $S_{Z}^{\rm B}$, $S_{X}^{\rm C}$, and $S_{Z}^{\rm D}$ are alternately performed to detect bit-flip or phase-flip errors in both the static and dynamical qubits. Each measurement result is compared to the result of the same stabilizer type from four rounds earlier, with the first four rounds compared to the encoding value. Errors during this process manifest as changes in the measurement outcomes, and the corresponding data are post-selected out in error detection. We present the preservation results for the encoded states $|-,0\rangle$ and $|-,1\rangle$ in Fig. 2e, showing the measured expectation values of the operator $X_{\rm s}Z_{\rm d}$. The logical error rates ($\varepsilon_{\rm exp}$) for these states, extracted by fitting the exponential decay of the expectation values, are reduced from 14.7% and 14.3% to 2.6% and 2.5%, respectively. This improvement corresponds to an increase in the logical lifetime $T_{\rm L}=\tau_{\rm exp}/2\varepsilon_{\rm exp}$ from 3.1 $\mu$s and 3.2 $\mu$s to 17.7 $\mu$s and 18.4 $\mu$[17], respectively. Here, $\tau_{\rm exp}=920$ ns is the duration of a stabilizer round. Numerical simulation results using the open-source package Stim [59], based on monitored gate and measurement errors, are also shown for comparison. The larger error rates observed in experiments may arise from leakage-induced correlated errors [60, 61, 62, 54, 63, 64, 58]. We measure an average leakage rate of 0.18% per round on the physical qubits using three-state readout at the end of the circuit (see Supplementary Information IV). In both the experimental and simulation data, round-dependent oscillations are observed. Specifically, higher expectation values occur every four rounds, corresponding to the completion of a full error detection cycle. Beyond error detection, error correction can be applied to the static qubit alone (see Supplementary Information VI). However, the effectiveness of error correction diminishes for the two-qubit states due to the small code distance of the dynamical qubit and the correlation between errors in the two logical qubits.
We demonstrate universal logical gates on the dynamical qubit, including FT Pauli gates and nFT arbitrary rotation gates. For logical gates on the static qubit, see Supplementary Information Section V and reference [12]. The Pauli gates are implemented transversally using parallel single-qubit gates on the corresponding data qubits, as defined by the logical operators. We insert the Pauli gates {$I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, $I_{\rm s}\otimes Z_{\rm d}$} between the stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as illustrated in Fig 3a. The final states, after applying the gates and conducting four stabilizer rounds, are characterized using LQST. For each encoded Pauli state of the dynamical qubit, the static qubit traverses four states: $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}$. The average state infidelities after applying the gates $I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, and $I_{\rm s}\otimes Z_{\rm d}$ are $0.255\pm 0.011$, $0.263\pm 0.013$, and $0.256\pm 0.011$, respectively. These values are close to the infidelity $0.253\pm 0.011$ observed for the identity operation (i.e., no gates inserted), indicating that the FT Pauli gates induce negligible errors.
The construction of arbitrary rotation gates is facilitated by the use of ancilla qubits. For logical rotations around the $Z_{\rm d}$-axis ($R_{Z}^{\varphi}$), a physical gate $R_{z}^{\varphi}$ is inserted on an ancilla qubit and placed between four physical CNOT gates acting on $D_{7}$ and $D_{8}$, as illustrated at the top of Fig. 3b. The entire circuit is then inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. Similarly, logical rotations around the $X_{\rm d}$-axis ($R_{X}^{\theta}$) are implemented using an $R_{z}^{\theta}$ gate on an ancilla qubit. The corresponding circuit is inserted between stabilizers $S_{X}^{\rm C}$ and $S_{Z}^{\rm D}$. We sweep the rotation angles $\varphi$ and $\theta$ and measure the dynamical qubit in three bases: $X_{\rm d}$, $Y_{\rm d}$, and $Z_{\rm d}$. The measurement results are shown in Fig. 3b, with filled (open) symbols representing data with (without) error detection. With error detection enabled, the lower amplitude in the $Y_{\rm d}$-direction is primarily due to the nFT nature of the logical $Y_{\rm d}$ measurement. Without error detection, the amplitude in the $Y_{\rm d}$-direction is lower than that in the $Z_{\rm d}$-direction for the $R_{X}^{\theta}$ rotation due to more phase-flip errors propagating through the nFT $R_{X}^{\theta}$ gate.
Figure 4: Logical CNOT gate between the dynamical and static qubits. a, The CNOT circuit (top) is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ in the overall circuit (bottom). For Bell state generation, the logical qubits are initialized to $|+,0\rangle$ in the encoding process. For LQPT, 16 linearly independent logical cardinal states: $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are encoded. b, Error-detected logical density matrix of the generated logical Bell state, obtained with four stabilizer rounds. The logical state fidelity is 75.9% with error detection and 38.8% without error detection. c, Extracted LPTM $R_{\rm exp}$ of the CNOT gate, with a process fidelity $F_{\rm p}=80.2\%$.
Finally, we show that the dynamical and static qubits can be entangled using a logical CNOT gate. The specific CNOT circuit depends on the preceding stabilizer type. In our implementation, the CNOT circuit is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. This circuit involves 14 entangling gates acting on a subgroup of 3 data qubits and 2 ancilla qubits, as illustrated in Fig. 4a. The circuit is derived from using the static qubit operator $Z_{\rm s}=Z_{1}Z_{4}Z_{7}$ (which is equivalent to $Z_{\rm s}$ in Eq. 1 up to stabilizer $S_{Z}^{\rm D}$) to control the dynamical qubit operator $X_{\rm d}=X_{1}X_{7}$ after round $S_{Z}^{\rm B}$ (see Supplementary Information Section I for theoretical derivation). We note that the logical CNOT gate we implement is nFT because it allows errors to propagate between the 3 involved data qubits. The circuit is intentionally designed to involve only 3 data qubits to minimize error propagation.
Applying the logical CNOT gate to the encoded state $|+,0\rangle$ yields a logical Bell state $\frac{1}{\sqrt{2}}(|0,0\rangle+|1,1\rangle)$. The error-detected logical density matrix $\rho_{\rm B}$ is depicted in Fig. 4b, with a fidelity of 75.9%. Without error detection, the corresponding fidelity is only 38.8%. We benchmark the logical CNOT gate fidelity using LQPT. 16 linearly independent two-qubit logical states $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are prepared in the encoding process. For each initial state, two-qubit LQST is performed at the end of the circuit. Due to the susceptibility of LQPT to state preparation and measurement (SPAM) errors, we characterize the input states using the same circuit without inserting the logical CNOT gate [65]. The logical Pauli transfer matrix (LPTM) $R_{\rm exp}$, which linearly matches the input states to output states, is calculated using the 16 input-output pairs (see Methods). The resulting LPTM with error detection is shown in Fig. 4b. The process fidelity is calculated by comparing $R_{\rm exp}$ with the ideal transfer matrix $R_{\rm ideal}$, yielding $\mathcal{F}_{p}={\rm Tr}(R_{\rm exp}\cdot R_{\rm ideal})=80.2\%$. This corresponds to a gate fidelity $\mathcal{F}_{g}=(\mathcal{F}_{p}/E+1)/(E+1)$ of 84.1%, where $E=4$ is the Hilbert space dimension. The fidelity calculation based on LQPT suffers from the low fidelity of encoded $|{-}i\rangle$ state on either qubit. We can also estimate the logical CNOT gate fidelity by dividing the average output state fidelity by the average input state fidelity without $|{-}i\rangle$, indicating a gate fidelity of 93.6%.
III Discussion and outlook
In this work, we implement the FBS code on a superconducting quantum processor, encoding a dynamical logical qubit in addition to a conventional static logical qubit. We experimentally demonstrate universal two-qubit logical operations integrated with repetitive quantum error detection, including encoding, measurement, arbitrary single-qubit rotations, and a CNOT gate. Some of these logical operations are implemented in a FT manner and outperform their nFT counterparts. Our first demonstration of universal logical operations with a dynamical logical qubit highlights the significant role played by Floquet codes in achieving resource-efficient FT quantum computation.
There are several challenges that remain for future work: (1) Despite improvements in code rate using the FBS code, our experiments show that the dynamical logical qubit is more susceptible to errors, requiring a larger code distance for effective protection. Specifically, the code distance for the dynamical qubit is approximately $d/2$ when the FBS code is implemented on a $d\times d$ square lattice [42]. Further research is needed to scale up the FBS code to larger lattices to enable effective error correction for the dynamical qubit. (2) Implementing FT entanglement and non-Clifford gates with the dynamical logical qubit is a critical next step. It has been shown that standard methods used in surface code, such as lattice surgery [24] and braiding twist defects [38], can be adapted to Floquet codes. A recent theoretical work has explored FT logical operations using a generalized Floquet code in a honeycomb lattice, demonstrating that entanglement operations can be achieved by simply reordering the measurement schedule [66]. This suggests that Floquet codes may offer a simpler and more natural pathway to FT quantum computing.
IV Methods
IV.1 Logical operator definition in the FBS code
The FBS code is derived from its parent subsystem code, the $[[9,1,3]]$ BS code, which employs a $3\times 3$ square lattice of data qubits to encode one logical qubit. The gauge group $G$ of the BS code is generated by the following $12$ weight-$2$ gauge operators (refer to Fig. 1a for qubit numbering):
$\langle$ $Z_{1}Z_{2},Z_{2}Z_{3},Z_{4}Z_{5},Z_{5}Z_{6},Z_{7}Z_{8},Z_{8}Z_{9},$
$X_{1}X_{4},X_{2}X_{5},X_{3}X_{6},X_{4}X_{7},X_{5}X_{8},X_{6}X_{9}\rangle.$
By identifying the operators in $G$ that commute with all elements in $G$, we obtain $4$ weight-$6$ stabilizers:
$S_{X}^{\rm A}$ $=X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},\quad S_{Z}^{\rm B}=Z_{2}Z_{3}Z_{5}Z_{6}Z_{8}Z_{9},$
$S_{X}^{\rm C}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6},\quad S_{Z}^{\rm D}=Z_{1}Z_{2}Z_{4}Z_{5}Z_{7}Z_{8}.$
The values of the above stabilizers can be obtained by measuring all $12$ gauge operators and multiplying the corresponding measurement outcomes. By identifying the operators that commute with all elements in $G$ but are not themselves in $G$, we obtain the logical operators of the static logical qubit, as defined in Eq. 1.
The FBS code is constructed by measuring the gauge operators in a period-$4$ schedule (see Fig. 1b), thereby encoding an additional dynamical logical qubit. Up to a ±$\pm 1$ sign, the logical operator of the dynamical qubit evolves according to the period-$4$ schedule, as shown in Table 1. In the remainder of this section, we omit the superscript “Q” if it is clear from the context.
Table 1: Expressions for unsigned dynamical logical operators.
Round Q A B C D
$X_{\rm d}^{\rm Q}$ $X_{1}X_{4}$ $X_{1}X_{7}$ $X_{4}X_{7}$ $X_{3}X_{9}$
$Z_{\rm d}^{\rm Q}$ $Z_{1}Z_{3}$ $Z_{7}Z_{8}$ $Z_{7}Z_{9}$ $Z_{8}Z_{9}$
$Y_{\rm d}^{\rm Q}$ $Y_{1}Z_{3}X_{4}$ $X_{1}Y_{7}Z_{8}$ $X_{4}Y_{7}Z_{9}$ $X_{3}Z_{8}Y_{9}$
For Pauli operators $P\in\{X,Y,Z\}$, the sign at round $r$ is given by formula $\Gamma_{P}^{(r)}=\prod_{i=1}^{r}\gamma_{P}^{(i)}$, where $\gamma_{P}^{(i)}$ is defined in Table 2. Combining all of the above, the logical $P$ operator of the dynamical qubit at round $r$ is given by $P_{\rm d}^{(r)}=\Gamma_{P}^{(r)}P_{\rm d}^{\rm Q},$ where ${\rm Q}={\rm A},{\rm B},{\rm C},{\rm D}$ for $r\equiv 1,2,3,0\pmod{4}$, respectively.
Table 2: Definition of $\gamma^{(i)}_{X},\gamma^{(i)}_{Z},\gamma^{(i)}_{Y}$. At round $i$, $x_{uv}^{(i)},z_{uv}^{(i)}\in\{\pm 1\}$ denotes the ancilla measurement outcome of weight-$2$ Pauli operators $X_{u}X_{v}$ and $Z_{u}Z_{v}$. $S_{X}^{\rm A},S_{Z}^{\rm B},S_{X}^{\rm C},S_{Z}^{\rm D}\in\{\pm 1\}$ refer to the initial values of the stabilizers determined by the encoding process.
$i~{}\mathrm{mod}~{}4$ 1 2 3 0
$\gamma_{X}^{(i)}$ $x_{69}^{(i)}x_{25}^{(i)}S_{X}^{\rm C}$ $x_{47}^{(i-1)}$ $x_{14}^{(i)}$ $x_{36}^{(i-1)}x_{58}^{(i-1)}S_{X}^{\rm A}$
$\gamma_{Z}^{(i)}$ $z_{12}^{(i-1)}z_{56}^{(i-1)}S_{Z}^{\rm B}$ $z_{23}^{(i)}z_{45}^{(i)}S_{Z}^{\rm D}$ $z_{89}^{(i-1)}$ $z_{78}^{(i)}$
$\gamma_{Y}^{(i)}$ $\gamma_{X}^{(i)}\gamma_{Z}^{(i)}$
IV.2 Logical operator evaluation using data-qubit measurement outcomes
For the static qubit, the logical operator is calculated from the measurement outcomes of data qubits during logical measurement. The values of the static logical operators are evaluated by multiplying the measurement outcomes of the corresponding data qubits, i.e., $X_{\rm s}=x_{4}x_{5}x_{6},\quad Z_{\rm s}=z_{2}z_{5}z_{8},\quad Y_{\rm s}=x_{4}x_{6}y_{5}z_{2}z_{8},$ where $x_{i},y_{i},z_{i}\in\{\pm 1\}$ denote the $X$, $Y$, and $Z$ measurement outcomes of data qubit $D_{i}$, respectively.
Assume the logical measurement is performed after $r$ rounds of stabilizer measurement. For Pauli operator $P\in\{X,Y,Z\}$, the value of the dynamical logical operator $P^{\rm(r)}_{\rm d}=\Gamma_{P}^{(r)}P_{\rm d}$ is the product of two parts: the value of its unsigned operator $P_{\rm d}$ and the value of its sign $\Gamma_{P}^{(r)}$ at round $r$. Similar to the static logical operators, the value of $P_{\rm d}$ is obtained by multiplying the measurement outcomes of the corresponding data qubits (see Table 1). The value of $\Gamma_{P}^{(r)}$ is calculated as $\Gamma_{P}^{(r)}=\prod_{i=1}^{r}\gamma_{P}^{(i)}$, where $\gamma_{P}^{(i)}$ is determined by the ancilla measurement outcomes during encoding and stabilizer measurements (see Table 2). Due to the presence of $\Gamma_{P}^{(r)}$, the evaluation of logical operators will change if we apply the logical rotation and the CNOT gate described in the main text (see Supplementary Information Section I for theoretical details).
IV.3 Two-qubit Logical Quantum State Tomography
The construction of the logical density matrix $\rho_{\rm L}$ in the $Z$ basis follows a procedure similar to that used for a physical density matrix. For a physical density matrix, single-qubit gates $\{I,R_{y}^{\pi/2},R_{x}^{\pi/2}\}$ are applied to the physical qubits before performing projective measurements, enabling measurements in the $X$, $Y$, and $Z$ directions. However, for logical states, these rotation operations are unnecessary. Instead, we directly perform logical measurements in the $X$, $Y$, and $Z$ bases. For a two-qubit logical state, we measure the state in 9 logical bases $\{Z,X,Y\}^{\otimes 2}$. For each basis, we obtain an array $[p_{--},p_{-+},p_{+-},p_{++}]$, where $p_{-+}$ denotes the probability of measuring the static qubit as $-1$ and the dynamical qubit as $+1$.
We optimize the logical density matrix $\rho_{\rm L}$ using the convex optimization package cvxpy [67], constraining the density matrix to be Hermitian, unit-trace and positive semi-definite. The fidelity of the reconstruted state relative to a target pure state is calculated as ${\rm Tr}(\rho_{\rm L}\rho_{\rm ideal})$.
IV.4 Two-qubit Logical Quantum Process tomography
LQPT is built upon the results of LQST. Each two-qubit logical density matrix $\rho_{\rm L}$ can be expressed in the two-qubit Pauli basis as a vector $\vec{p}^{\mathrm{T}}=[p_{00},p_{01},...,p_{32},p_{33}]$, where
$p_{ij}=\frac{{\rm Tr}(\rho_{\rm L}\sigma_{i}\otimes\sigma_{j})}{{\rm Tr}(\rho_{\rm L})},\text{and}~{}\sigma_{i},\sigma_{j}\in\{I,X,Y,Z\}.$
The LPTM $R$, a $16\times 16$ matrix, maps an input state vector $\vec{p}$ to an output state vector $\vec{p^{\prime}}=R\vec{p}$ [68]. To reconstruct the LPTM, we employ convex optimization to find an optimal LPTM that satisfies all the constraints of a physical quantum channel, using an overcomplete set of input states.
Acknowledgements
We thank Bujiao Wu for valuable discussions. This work was supported by the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034), the National Natural Science Foundation of China (12174178, 12374474, 123b2071).
Author contributions
X.D.S. developed the FPGA program for custom electronics, collected and analyzed the data. L.C.L. conducted the theoretical analysis, supervised by X.M.S. Z.Y.W. and J.C. calibrated the quantum processor. J.C. conceived the experiments. Y.P.Z. and D.P.Y. supervised the project. All authors contributed to the experimental setup, discussions of the results and writing of the manuscript.
Competing interests
The authors declare no competing interests.
Data availability
The data that support the plots within this paper and other findings of this study are available from the corresponding authors upon reasonable request.