### v4

# Logical operations with a dynamical qubit in Floquet-Bacon-Shor code

Abstract
Quantum error correction (QEC) protects quantum systems against inevitable noises and control inaccuracies, providing a pathway towards fault-tolerant (FT) quantum computation. Stabilizer codes, including surface code and color code, have long been the focus of research and have seen significant experimental progress in recent years. Recently proposed time-dynamical QEC, including Floquet codes and generalized time-dynamical code implementations, opens up new opportunities for FT quantum computation. By employing a periodic schedule of low-weight parity checks, Floquet codes can generate additional dynamical logical qubits, offering enhanced error correction capabilities and potentially higher code performance. Here, we experimentally implement the Floquet-Bacon-Shor code on a superconducting quantum processor. We encode a dynamical logical qubit within a $3\times 3$ lattice of data qubits, alongside a conventional static logical qubit. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using repeated error detection. We showcase universal single-qubit logical gates on the dynamical qubit. Furthermore, by implementing a logical CNOT gate, we entangle the dynamical and static logical qubits, generating an error-detected logical Bell state with a fidelity of 75.9%. Our results highlight the potential of Floquet codes for resource-efficient FT quantum computation.
Figure 1: Implementation of the FBS code. (a) Schematic of the quantum device for the FBS code, consisting of 9 data qubits (white circles labeled as $D_{i}$, where i = 1 to 9) and 12 ancilla qubits (orange circles for $Z$ ancillas and blue circles for $X$ ancillas). The four central qubits are alternately used as $X$ and $Z$ ancillas in different stabilizer rounds. Extra ancilla qubits (grey circles) are used for circuit compression and the construction of logical CNOT gate. Each pair of qubits is connected by a tunable coupler (H shape). (b) Illustration of stabilizer measurements for the BS code (left) and the FBS code (right). The FBS code features period-4 stabilizer measurements, during which the logical operators $X_{\rm d}$ (purple lines) and $Z_{\rm d}$ (green lines) for the dynamical qubit evolve across different rounds. In contrast, the logical operators $X_{\rm s}$ and $Z_{\rm s}$ for the static qubit remain fixed and are identical in both codes. Weight-2 Pauli measurements are used to check the parity of the corresponding data qubits in the $X$ (blue) and $Z$ (orange) bases. (c) Cumulative distributions of simultaneous errors for single-qubit gates, CZ gates, ancilla qubit readout, and data qubit idle (with dynamical decoupling) during ancilla readout are depicted. The median error rates are 0.07%, 0.97%, 1.58%, and 1.43%, respectively, as indicated by the vertical lines.
Quantum computers hold potential for solving problems intractable for classical systems [1, 2, 3]. However, their practical realization faces the fundamental challenge of quantum information fragility. To address this challenge, quantum error correction (QEC) has been established as a critical framework [4, 5, 6]. By redundantly encoding logical qubits into larger systems with many physical qubits, QEC detects and corrects errors through repeated stabilizer measurements. QEC codes, including surface code [7], color code [8], Bacon-Shor (BS) code [9], and quantum LDPC codes [10], have made significant experimental progress across various systems in recent years [11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Particularly in logical state preservation, exponential error suppression has been achieved by scaling a surface-code logical qubit in superconducting circuits [21]. Beyond state preservation, logical operations—such as arbitrary single-qubit gates and entangling gates between logical qubits—are essential for fault-tolerant (FT) quantum computing [22]. However, implementing these operations is generally challenging, especially in physical systems with local connectivity [23, 24, 25].
Recently proposed time-dynamical QEC, including Floquet codes [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] and generalized time-dynamical code implementations [50, 51, 52, 53, 54, 55], opens up new opportunities for FT quantum computation by transcending the usual paradigm of stabilizer codes or subsystem codes, where logical information is encoded in a fixed subspace. The Floquet codes hold the advantage of requiring only low-weight Pauli measurements, typically of weight-2 [26], which are naturally compatible with physical architectures supporting native two-qubit measurements [56, 57]. Moreover, by periodically encoding logical information over time, Floquet codes enable additional dynamical logical qubits compared to their parent codes, offering greater flexibility and potentially higher code performance [39].
In this work, we implement the Floquet-Bacon-Shor (FBS) code [41] on a superconducting quantum processor. As a minimal, distance-2 implementation of a Floquet code, it requires only square-lattice connectivity and is thus readily implementable on contemporary hardware. Compared to the conventional BS code, the FBS code additionally encodes one dynamical logical qubit, enabling two-qubit logical operations. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using multi-round stabilizer measurements. We showcase FT Pauli gates and non-fault-tolerant (nFT) arbitrary rotation gates on the dynamical qubit. Furthermore, we achieve a logical Bell state with a fidelity of $75.9\%$ by entangling the dynamical and static logical qubits via a logical CNOT gate, with the gate fidelity assessed at $84.1\%$ through logical quantum process tomography (LQPT). Our results highlight the potential of Floquet codes for scalable and resource-efficient QEC.
Figure 2: Encoding, stabilization, and measurement of two-qubit logical states in the FBS code. (a) Simplified circuit for encoding, stabilizing and measuring two-qubit logical states. All physical qubits are initialized to the ground state before the circuit. Period-4 stabilizer measurements are alternately performed following the encoding circuit. (b) Encoding circuits for logical Pauli states $|+,0\rangle$ and $|1,0\rangle$. BS stabilizer measurements are included in the latter case, with the dynamical qubit state post-selected to be $|0\rangle$. The notation $D_{147}$ denotes the three data qubits $D_{1}$, $D_{4}$, and $D_{7}$, and so on. The CNOT symbol denotes transversal CNOT gates between data and ancilla qubit groups, subject to connectivity constraints. (c) Example logical measurement circuits after stabilizer $S_{Z}^{\rm D}$: FT ZZZZ measurement (left) and nFT XZXZ measurement (right). Logical operators are indicated within the dashed-line boxes. Measurement outcomes of data qubits are used to calculate the values of the $X$ (blue) and $Z$ (orange) stabilizers, enabling error detection. (d) LQST fidelities of all 36 two-qubit Pauli states. The 16 states $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}^{\otimes 2}$ are fault-tolerantly prepared and exhibit higher fidelity, as shown on the left side (white background). The blue portions indicate the raw fidelities (i.e., without using syndrome data), while the green portions highlight the fidelity improvement by error detection (i.e., with syndrome data post-selected). (e) Experimentally measured (filled symbols) and simulated (open symbols) expectation values of $X_{\rm s}Z_{\rm d}$ operator versus stabilizer round, for the encoded states $|-,1\rangle$ (left) and $|-,0\rangle$ (right). The solid and dashed lines represent exponential fits. Each stabilizer round lasts 920 ns, comprising 720 ns for ancilla readout and 200 ns for single-qubit and two-qubit gates. The cyan and green colors correspond to the raw data and error-detected data, respectively. The dashed grey lines indicate the average error rates of physical two-qubit states. Due to the exponential decrease in the retained data during error detection, the analysis is limited to 16 rounds, with a total of 100,000 experimental shots. The insets show the error detection probabilities for the weight-6 stabilizers over 24 measurement rounds, with the final round calculated from the logical measurements.
The FBS code we demonstrate is a generalization of the $[[9,1,3]]$ BS code [9], which encodes one logical qubit in a $3\times 3$ square lattice, as illustrated in Fig. 1(a). Each labeled vertex represents a data qubit, and other colored vertices represent ancilla qubits for weight-2 Pauli measurements. The static logical qubit is defined within a fixed subsystem by the operators:
$X_{\rm s}=X_{4}X_{5}X_{6},\quad Z_{\rm s}=Z_{2}Z_{5}Z_{8},\quad Y_{\rm s}=iX_{\rm s}Z_{\rm s}.$ (1)
The BS code is stabilized by weight-2 Pauli measurements [9], whose products yield four weight-6 stabilizers $S_{X}^{\text{A}}$, $S_{Z}^{\text{B}}$, $S_{X}^{\text{C}}$, and $S_{Z}^{\text{D}}$, enabling the detection and correction of any single-qubit error. Since these Pauli measurements do not always commute, the Hilbert space of the 9 physical qubits consists of 4 degrees of freedom fixed by the stabilizers, 1 degree for logical information, and 4 additional degrees corresponding to the gauge space, which does not encode logical information. By imposing a period-4 ordering on the weight-2 Pauli measurements, the gauge space dynamically encodes an additional logical qubit, thereby forming the FBS code. The logical Pauli operators $X_{\text{d}}$ and $Z_{\text{d}}$ of the dynamically encoded qubit evolve periodically during the stabilizer measurements, as indicated by the purple and green lines in Fig. 1(b). Throughout this paper, the initial operator definitions are consistent with stabilizer round A. Although the $3\times 3$ system size is insufficient for error correction of the dynamical qubit, the stabilizers extracted from the measurements are sufficient to detect any single-qubit error. Within the framework of error detection, an operation is considered FT if any single-qubit fault produces a non-trivial syndrome and is thereby detectable [58].
We implement the FBS code on a 21-qubit subsystem of a 66-transmon superconducting processor [59] with 110 tunable couplers [60]. The $N_{d}=9$ data qubits, labeled as $D_{i}$ ($i=1$ to 9), form a $3\times 3$ array and are interlaced with $X$ and $Z$ ancilla qubits for stabilizer measurements. The cumulative distributions of simultaneous gate errors and readout errors [61] are illustrated in Fig. 1(c). The experimental setup, device parameters and calibration methods are detailed in the Supplementary Material [62].
Figure 3: Logical gates on the dynamical qubit. (a) FT Pauli gates on the dynamical qubit. Transversal Pauli gates are applied to the data qubits and inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as shown at the top. The final state is characterized using LQST. The state infidelities with error detection after applying the Pauli gates on different encoded states are shown below the circuit. For each dynamical qubit state, the infidelity is averaged over 4 static qubit states $|0\rangle$, $|1\rangle$, $|+\rangle$, $|$-$\rangle$}. Error bars correspond to 95% confidence intervals. (b) Logical rotation gates on the dynamical qubit around the $Z_{\rm d}$-axis (left) and $X_{\rm d}$-axis (right). The measurement expectation values of the operators $X_{\rm d}$ (cyan), $Y_{\rm d}$ (green) and $Z_{\rm d}$ (orange) are shown as functions of rotation angles $\varphi$ and $\theta$. The gate circuits (shown at the top) for rotations around the $Z_{\rm d}$-axis ($X_{\rm d}$-axis) are inserted between stabilizers $S_{Z}^{\rm B}$ ($S_{X}^{\rm C}$) and $S_{X}^{\rm C}$ ($S_{Z}^{\rm D}$). The filled (open) symbols represent data with (without) error detection in the four stabilizer rounds and the final logical measurement. The dashed curves correspond to trigonometric fits.
[…]
We perform period-4 stabilizer measurements to preserve the encoded two-qubit logical states. In each stabilizer round, 4 weight-2 Pauli measurements are conducted. Each measurement result is compared to the result of the same stabilizer type from four rounds earlier, with the first four rounds compared to the initial encoding values. The preservation results for the encoded states $|-,0\rangle$ and $|-,1\rangle$ are presented in Fig. 2(e), showing the measured expectation values of the operator $X_{\rm s}Z_{\rm d}$. With error detection, the logical error rates per round ($\varepsilon_{\rm exp}$), extracted by fitting the exponential decay of the expectation values, are reduced from 14.7% and 14.3% to 2.6% and 2.5%, respectively. The error rates are compared to those of physical two-qubit states (dashed grey lines), with $\overline{\varepsilon}_{\rm phy}=1-\exp(-\tau_{\rm exp}/\overline{T_{2}^{\rm echo}})/2-\exp(-\tau_{\rm exp}/\overline{T_{1}})/2=4.4\%$. Here, $\tau_{\rm exp}=920$ ns is the duration of a stabilizer round, $\overline{T_{2}^{\rm echo}}=11.7~\mu s$ is the average spin-echo dephasing time, and $\overline{T_{1}}=77.1~\mu s$ is the average energy relaxation time of the physical qubits. Numerical simulation results based on monitored gate and readout errors are also shown for comparison [65]. The larger error rates observed in experiments may arise from leakage-induced correlated errors [62, 66, 67, 68, 69, 70, 71, 64]. In both the experimental and simulation data, round-dependent oscillations are observed, with higher expectation values occur every four rounds, corresponding to the completion of a full error detection cycle.
Figure 4: Logical CNOT gate between the dynamical and static qubits. (a) The CNOT circuit (top) is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ in the overall circuit (bottom). For Bell state generation, the logical qubits are initialized to $|+,0\rangle$ during the encoding process. For LQPT, 16 linearly independent logical cardinal states: $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are encoded. (b) Error-detected logical density matrix of the generated logical Bell state, obtained with four stabilizer rounds. The logical state fidelity is 75.9% with error detection and 38.8% without error detection. The wireframes indicate the ideal values. (c) Extracted LPTM $R_{\rm exp}$ of the CNOT gate, with a process fidelity $F_{\rm p}=80.2\%$.
We demonstrate universal logical gates on the dynamical qubit, including FT Pauli gates and nFT arbitrary rotation gates. For logical gates on the static qubit, refer to the Supplementary Material [62] and reference [11]. Pauli gates are implemented transversally via single-qubit gates on corresponding data qubits, inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ (Fig. 3(a)). LQST characterization reveals that the average state infidelities after applying the gates $I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, and $I_{\rm s}\otimes Z_{\rm d}$ are $0.255\pm 0.011$, $0.263\pm 0.013$, and $0.256\pm 0.011$, respectively. These values closely match the state infidelity without gate inserted ($0.253\pm 0.011$), confirming that the FT Pauli gates induce negligible errors.
Arbitrary logical rotations are constructed using ancilla qubits. The $R_{Z}^{\varphi}$ gate is realized by inserting a physical $R_{z}^{\varphi}$ gate on an ancilla between four CNOTs on $D_{7}$ and $D_{8}$, placed between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as illustrated at the top of Fig. 3(b). $R_{X}^{\theta}$ is constructed in a similar way and is inserted between $S_{X}^{\rm C}$ and $S_{Z}^{\rm D}$. Sweeping angles $\varphi$ and $\theta$ and measuring $X_{\rm d}$, $Y_{\rm d}$, and $Z_{\rm d}$ yields the results in Fig. 3(b), where filled (open) symbols denote data with (without) error detection. The reduced $Y_{\rm d}$ amplitude under error detection results from its nFT measurement. Without error detection, $R_{X}^{\theta}$ gate exhibits a further suppressed $Y_{\rm d}$ amplitude compared to $Z_{\rm d}$, a consequence of enhanced phase-flip error propagation through this nFT gate. Data post-processing is used to avoid real-time feedback [62].
Finally, we demonstrate entanglement between the dynamical and static qubits using a logical CNOT gate. The gate circuit, inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, involves 14 entangling gates acting on 3 data qubits and 2 ancilla qubits, as illustrated in Fig. 4(a). The circuit is derived by using the static qubit operator $Z_{\rm s}=Z_{1}Z_{4}Z_{7}$ (equivalent to $Z_{\rm s}$ in Eq. 1) to control the dynamical qubit operator $X_{\rm d}=X_{1}X_{7}$ after stabilizer $S_{Z}^{\rm B}$ [62]. Although the logical CNOT gate is nFT due to error propagation among the 3 data qubits, the circuit is designed to minimize such effects.
Applying the logical CNOT gate to $|+,0\rangle$ yields a logical Bell state $\frac{1}{\sqrt{2}}(|0,0\rangle+|1,1\rangle)$. The error-detected logical density matrix $\rho_{\rm B}$ is depicted in Fig. 4(b), with a fidelity of 75.9%. We benchmark the gate fidelity using LQPT with 16 input states $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$. For each input, two-qubit LQST is performed at the end of the circuit. To account for state preparation and measurement (SPAM) errors, we characterize the input states separately using the same circuit without the CNOT gate [23]. The resulting logical Pauli transfer matrix (LPTM) $R_{\rm exp}$ (Fig. 4(c)) is constructed from the 16 input-output pairs [62]. Comparing it with the ideal matrix $R_{\rm ideal}$ gives a process fidelity of $\mathcal{F}_{p}={\rm Tr}(R_{\rm exp}\cdot R_{\rm ideal})=80.2\%$, which corresponds to a gate fidelity $\mathcal{F}_{g}=(\mathcal{F}_{p}/E+1)/(E+1)=84.1\%$ for the $E=4$-dimensional Hilbert space [72]. The LQPT fidelity suffers from the relatively low fidelity of encoded $|{-}i\rangle$ states. As an alternative estimate, we exclude these states and compute the ratio of the average output state fidelity to the average input state fidelity, which yields a gate fidelity of 93.6%.
In conclusion, we implement the FBS code on a superconducting quantum processor, encoding both a dynamical logical qubit and a conventional static logical qubit. We demonstrate universal logical operations on the dynamical qubit, integrated with repeated quantum error detection, and entangle the dynamical and static logical qubits by constructing a logical CNOT gate. Future work may scale Floquet codes to larger lattices for dynamical qubit error correction. However, the standard FBS code cannot provide a practical scaling due to lack of a threshold [41]. Techniques such as code concatenation [73, 74] and measurement skipping [75, 76] could be adapted to enable effective error suppression at larger code distances. Additionally, methods like lattice surgery [27] and braiding twist defects [37] may be integrated into Floquet codes to enable FT entanglement and non-Clifford gates with the dynamical logical qubit.
We thank Bujiao Wu for valuable discussion. This work was supported by the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034), the Innovation Program for Quantum Science and Technology (2021ZD0301703), the National Natural Science Foundation of China (62325210, 12174178, 12374474, 123b2071).
The data that support the findings of this article are openly available [77].